OpenStudy (anonymous):

Emma finished a small bag of her favorite snack that weighed zero point three seven five pound. Wendy ate one-sixth of a one-pound bag of the snack. Who ate more? Explain.

OpenStudy (anonymous):

compare the numbers \[0.375\] with \(\frac{1}{6}\)

OpenStudy (anonymous):

you can compare the fractions \[\frac{3}{8}\] and \[\frac{1}{6}\] or the decimals \[0.375\] and \[.166666...\]

OpenStudy (anonymous):

evidently \(\frac{1}{6}\) is the smaller of the two numbers